### v1

@@ 5.1 Algorithmic Complexity Determines Model Performance Patterns
Models are generally proficient at algorithm tags that require basic mathematical procedures and minimal compositional reasoning. As shown in Table 3, all evaluated models consistently achieve higher pass rates on tasks categorized under implementation, mathematics, prefix sum, sorting, and graph traversal—GPT-5 notably attains over $70\%$ accuracy on most of these tags. Such tasks primarily depend on recognizing familiar solution templates or leveraging procedural knowledge obtained from training. Performance noticeably declines for algorithms demanding deeper analytical reasoning or succinct proofs, such as greedy methods and graph theory, where even top proprietary models like GPT-5 drop to around $60\%$. The greatest difficulties arise in tasks that require on-the-spot creative observations, intricate state designs, or hierarchical invariants—particularly evident in dynamic programming (DP), segment trees (ST), and tree (TR) problems, where GPT-5’s pass rate sharply decreases to approximately $47\%$, $56\%$, and $38\%$, respectively. To address these weaknesses, future work could explore curriculum-driven fine-tuning (Huang et al., 2025) using carefully designed synthetic datasets of complex graph, tree, and DP problems, encouraging models to internalize the recurrence relations, hierarchical invariants, and compositional reasoning patterns crucial to solving these more challenging algorithmic tasks.
@@ 5.2 Reasoning Trace Analyses: Stronger Models Allocate Reasoning Tokens More Strategically
Figure 3: Reasoning Trace Analyses. We categorized eight reasoning behaviors and divide them into five groups: Analysis (Algorithm/Proof analysis and Complexity Analysis), Planning (Problem Restatement and Subgoal Setting), Exploration (Backtracking and Dead‑end recognition), Implementation (Pseudo implementation), Verification (Test Case Verification).
To better understand how thinking models solve challenging competitive coding problems, we conduct a detailed analysis on models’ reasoning traces. Inspired by prior work (Gandhi et al., 2025; Ahmad et al., 2025) on reasoning behavior analysis, we categorize models’ reasoning traces into eight behaviors and classify them into five groups as shown in Figure 3. Each trace is segmented into shorter chunks and annotated using GPT-OSS-120B. More details on the annotation prompt and implementation can be found in Appendix E.3.
GPT-OSS-120B increases exploration and analysis with problem difficulty, yet maintains stable exploration levels across reasoning budgets. In Figure 3(a), on more challenging problems, GPT-OSS-120B-High devotes significantly more effort to exploration—searching for viable solution paths—and deeper problem analysis, simultaneously reducing the tokens spent on initial planning and verification. This indicates that initial problem structuring behaviors are typically conducted early and not revisited extensively once a potential solution path is identified. Notably, even when provided with increased reasoning budgets (from low to high reasoning effort), as shown in Figure 3(b), GPT-OSS-120B strategically allocates extra tokens toward analysis, implementation, and verification, rather than further exploration. By maintaining stable exploration levels despite increased reasoning resources, the model mitigates excessive pivoting, a critical behavior that could otherwise lead to inefficient or incomplete reasoning traces, or “underthink” (Shojaee et al., 2025).
@@
Initial planning behaviors and subsequent verification steps play crucial roles in models producing correct solutions. Building upon this observation, we also investigate which reasoning behaviors distinguish correct from incorrect solutions. As illustrated in Figure 3(d), correct solutions exhibit increased planning behaviors, potentially explaining why exploration behaviors diminish—well-structured planning facilitates clearly defined solution paths, reducing the need for exploratory detours. Additionally, correct solutions engage in verification behaviors more frequently, ensuring adequate solution checks. This increased verification slightly reduces the need for extensive analysis, as models rely less on continuous reevaluation once confident in their solution correctness. Notably, correct solutions include more verification because targeted end-checks consolidate successful trajectories; however, stronger models rely less on explicit verification overall due to robust upfront analysis and planning, which internalize many checks and reduce the need for post-hoc verification.
Based on these insights, an important direction for future research is to optimize how models allocate reasoning effort across different cognitive behaviors. Specifically, exploring methods that strategically guide model attention toward structured initial planning and in-depth analysis—while carefully balancing exploration to prevent excessive pivoting—could substantially enhance both solution correctness and reasoning efficiency. Additionally, developing mechanisms that help models internalize verification during planning and analysis phases could reduce their dependence on explicit, post-hoc verification steps. Finally, creating automated techniques to dynamically adjust reasoning budgets and behaviors according to problem-specific factors like difficulty and algorithms may further boost the effectiveness of reasoning models in solving complex competitive programming tasks.
@@ 5.3 Error Patterns in Model-Generated Code Submissions
As models exhibit stronger reasoning capabilities, their solutions show substantial reductions in failure types of time limit, memory limit, and compilation errors. However, runtime errors, although somewhat reduced, do not experience as pronounced a decline, highlighting persistent challenges in edge-case handling and execution robustness. We hypothesize that one possible reason top-performing models still exhibit relatively high runtime error rates could be their tendency to pursue more aggressive and optimized coding patterns, such as employing custom data structures, in-place transformations, and pointer arithmetic. These advanced techniques, while algorithmically sound, might inherently increase the potential for execution faults [^6] For instance, a simple algorithm like summing elements of an array becomes significantly more complex when highly optimized for memory access patterns using techniques such as loop unrolling and pragma directives in C++ (e.g., #pragma omp simd, #pragma unroll)., especially in edge scenarios. Interestingly, GPT-OSS-20B displays compilation error rates comparable to weaker, non-reasoning-intensive models. We attribute this unexpected result to its cautious approach: the model often declines to generate solutions when it anticipates insufficient reasoning time, thereby triggering compilation-related failures. These findings highlight a limitation in the reinforcement learning approaches employed by current models (DeepSeek-AI et al., 2025; Yang et al., 2025b), which predominantly use solution correctness as the sole reward, neglecting efficiency and memory management. Future training techniques could incorporate fine-grained reward signals targeting these attributes, enabling models to optimize not only for correctness but also for reliable and efficient code execution.
@@ 6 Conclusion
In this work, we propose LiveOIBench, a comprehensive competitive coding benchmark featuring expert-curated OI-style tasks with detailed subtask rubrics, direct comparisons to human contestant performance, continuous updates with new Olympiad tasks to prevent contamination, and an offline evaluation system ensuring accessible and reproducible assessments. We extensively evaluate $32$ models including both proprietary and open-weight models. Our results highlight that proprietary models, particularly GPT-5, achieve impressive results but fall short of top human contestants, who typically place above the 90th percentile. Among open-weight models, GPT-OSS, Seed-OSS, and Qwen-3-32B demonstrate significant progress, with GPT-OSS-120B notably narrowing the performance gap to proprietary alternatives. Further analyses reveal that current models particularly struggle with advanced algorithmic tasks, such as dynamic programming. Additionally, our reasoning trace analysis indicates that robust model performance relies on strategically allocating exploratory and analytical reasoning behaviors. Lastly, we find stronger models reduce common failures yet persistently face runtime errors due to optimized coding techniques, suggesting refined training for efficiency and memory management. Moving forward, we envision leveraging this benchmark to further investigate inference-time scaling strategies and training methods, particularly for challenging reasoning tasks. By offering a rigorous, reproducible, and continuously updated evaluation benchmark, LiveOIBench aims to drive significant advancements in the reasoning and coding capabilities of LLMs.
@@ A.1 Task Collection
Contest Information Extraction: We developed a dedicated web crawler for each competition to extract task information directly from its official website. This includes task statements, test cases, reference and unofficial solutions, code attachments, time and memory limits, and detailed subtask specifications. We also parsed the contestant results pages and reformatted them into standardized CSV files. To do this, we copied the raw webpage content into Gemini‑2.5‑Pro and prompted it to generate CSVs with normalized headers. Each file captures contestant names, countries, total and per‑task scores, and awarded medals. After manual verification against the official data, we integrated the processed results into our contestant database. After integrating contestant results into our database, we determined medal thresholds as follows. For general contests, the Gold, Silver, and Bronze thresholds are defined by the lowest total scores among participants who received each respective medal. In contrast, for the USACO Bronze, Silver, and Gold contests, thresholds correspond to the minimum scores required to advance to the next competition level. In the case of the USACO Platinum contest, thresholds are based solely on the number of problems solved: solving exactly one problem earns a Bronze medal, two problems earn Silver, and solving more than two problems earns Gold.
@@ A.4 Codeforces Ratings Collection
Model Ratings: To benchmark model performance on our dataset, we calculate a corresponding Codeforces rating for each model on every contest. For each task, we present the full problem statement to the models and prompted them to generate code solutions. Using the provided subtask and test-case data, we compute the total score of each model’s solution. Once total scores were obtained, we derive Codeforces ratings for the models using the CodeElo formula (Quan et al., 2025) given below, where $m$ is the expected rank of a contestant (or model) with rating $r$, compared to $n$ contestants with known Codeforces ratings $r_{(i)}$.
@@
Competitions
Total
Tasks
Total
Contests
Avg.
Subtasks
Test
Cases/Task
Token
Count
Difficulty
@@
Table A4: Statistics of different competitions. USACO doesn’t provide subtasks information.
@@
Contest Contestants Medium Rating
@@ Appendix B Model Information
  • Proprietary LLMs: This category includes high-performing proprietary models such as Gemini-2.5 (Comanici et al., 2025), GPT-o3-Mini-High (OpenAI, 2025c), and GPT-4.1 (OpenAI, 2025a).
@@
  • Open-weight Thinking LLMs: These are openly available models that are empowered with inherent thinking or reasoning capabilities. This group includes Qwen3 (Yang et al., 2025b) and DeepSeek-R1 (DeepSeek-AI et al., 2025), as well as those distilled from DeepSeek-R1.
@@
  • Open-weight Non-Thinking LLMs: This category consists of openly available models that are not equipped with intrinsic thinking mechanisms. This includes DeepSeek Coder-V2 (DeepSeek-AI et al., 2024b), DeepSeek-V3 (DeepSeek-AI et al., 2024a), Qwen2.5 (Yang et al., 2024), Qwen2.5-Coder (Hui et al., 2024), Qwen3 (Yang et al., 2025a), Mistral (Jiang et al., 2023) and Llama-3 (Dubey et al., 2024).
@@
  • Refer to Table A7 and Table A8 for more details.
@@
Table A7: Model list of Non-Thinking LLMs with model providers
Non-Thinking LLMs Model Provider
GPT-4.1 (OpenAI, 2025a) OpenAI
Qwen2.5-72B (Yang et al., 2024) Alibaba
Qwen2.5-Coder-32B-Instruct (Hui et al., 2024) Alibaba
Qwen2.5-Coder-14B-Instruct (Hui et al., 2024) Alibaba
Qwen2.5-Coder-7B-Instruct (Hui et al., 2024) Alibaba
Mistral-Large-Instruct-2411 (Jiang et al., 2023) Mistral
Mistral-Small-3.1-24B-2503 (Jiang et al., 2023) Mistral
Llama-4-Scout (Meta AI, 2025) Meta
Llama-3.3-70B-Instruct (Dubey et al., 2024) Meta
Llama-3.1-8B-Instruct (Dubey et al., 2024) Meta
DeepSeek-V3 (DeepSeek-AI et al., 2024a) DeepSeek
DeepSeek-Coder-V2-Lite-Instruct (DeepSeek-AI et al., 2024b) DeepSeek
Codestral-22B-v0.1 (Mistral AI team, 2024) Mistral
@@
Table A8: Model list with categories, including model names, organizations, and reasoning budget
@@
Thinking LLMs Model Provider Reasoning Budget
GPT-5 (OpenAI, 2025b) OpenAI Medium
GPT-O3-Mini-High (OpenAI, 2025c) OpenAI High
GPT-OSS-120B-High (OpenAI et al., 2025) OpenAI High
GPT-OSS-20B-High (OpenAI et al., 2025) OpenAI High
GPT-OSS-120B (OpenAI et al., 2025) OpenAI Medium
GPT-OSS-20B (OpenAI et al., 2025) OpenAI Medium
SEED-OSS (ByteDance Seed Team, 2025) ByteDance Unlimited
Qwen3-32B (Yang et al., 2025a) Alibaba 38k
Qwen3-14B (Yang et al., 2025a) Alibaba 38k
QwQ-32B (Qwen Team, 2025) Alibaba 32K
Qwen3-30B (Yang et al., 2025a) Alibaba 38k
Qwen3-8B (Yang et al., 2025a) Alibaba 38k
Qwen3-4B (Yang et al., 2025a) Alibaba 38k
Gemini-2.5-Pro-exp-03-25 (Comanici et al., 2025) Google 64k
Gemini-2.5-Flash-preview-04-17 (Comanici et al., 2025) Google 64k
DeepSeek-R1-01-28 (DeepSeek-AI et al., 2025) DeepSeek 32k
DeepSeek-R1-Distill-Llama-70B (DeepSeek-AI et al., 2025) DeepSeek 32k
DeepSeek-R1-Distill-Qwen-32B (DeepSeek-AI et al., 2025) DeepSeek 32k
DeepSeek-R1-Distill-Qwen-14B (DeepSeek-AI et al., 2025) DeepSeek 32k
DeepSeek-R1-Distill-Llama-8B (DeepSeek-AI et al., 2025) DeepSeek 32k

### v2

@@ 5.1 Algorithmic Complexity Determines Model Performance Patterns
Models are generally proficient at algorithm tags that require basic mathematical procedures and minimal compositional reasoning. As shown in Table 3, all evaluated models consistently achieve higher pass rates on tasks categorized under implementation, mathematics, prefix sum, sorting, and graph traversal—GPT-5 notably attains over $70\%$ accuracy on most of these tags. Such tasks primarily depend on leveraging similar concepts or procedural knowledge obtained from training. Performance noticeably declines for algorithms demanding deeper analytical reasoning, such as greedy methods and graph theory, where even top proprietary models like GPT-5 drop to around $60\%$. The greatest difficulties arise in tasks that require on-the-spot creative observations, intricate state designs, or hierarchical invariants—particularly evident in dynamic programming (DP), segment trees (ST), and tree (TR) problems, where GPT-5’s pass rate sharply decreases to approximately $47\%$, $56\%$, and $38\%$, respectively. To address these weaknesses, future work could explore curriculum-driven fine-tuning (huang2025rzeroselfevolvingreasoningllm) using carefully designed synthetic datasets of complex graph, tree, and DP problems, encouraging models to internalize the hierarchical invariants and compositional reasoning patterns crucial to solving these more challenging algorithmic tasks.
@@ 5.2 Reasoning Trace Analyses: Stronger Models Allocate Reasoning Tokens More Strategically
Figure 3: Reasoning Trace Analyses. We categorize eight reasoning behaviors and divide them into five groups: Analysis (Algorithm/Proof analysis and Complexity Analysis), Planning (Problem Restatement and Subgoal Setting), Exploration (Backtracking and Dead‑end recognition), Implementation (Pseudo implementation), Verification (Test Case Verification).
To better understand how thinking models solve challenging competitive coding problems, we conduct a detailed analysis on models’ reasoning traces. Inspired by prior work (gandhi2025cognitivebehaviorsenableselfimproving; ahmad2025opencodereasoningadvancingdatadistillation) on reasoning behavior analysis, we categorize models’ reasoning traces into eight behaviors and classify them into five groups as shown in Figure 3. Each trace is segmented into shorter chunks and annotated using GPT-OSS-120B. More details on the annotation prompt and implementation can be found in Appendix E.8.
GPT-OSS-120B increases exploration and analysis with problem difficulty, yet maintains stable exploration levels across reasoning budgets. In Figure 3(a), on more challenging problems, GPT-OSS-120B-High devotes significantly more effort to exploration and deeper problem analysis, while it spends less tokens on initial planning and verification. This indicates that initial problem structuring behaviors are typically conducted early and not revisited extensively once a potential solution path is identified. Notably, even when provided with increased reasoning budgets (from low to high reasoning effort), as shown in Figure 3(b), GPT-OSS-120B strategically allocates extra tokens toward analysis, implementation, and verification, rather than further exploration. By maintaining stable exploration levels despite increased reasoning resources, the model mitigates excessive pivoting, a critical behavior that could lead to inefficient reasoning traces, or “underthink” (DBLP:journals/corr/abs-2506-06941).
@@
Initial planning behaviors and subsequent verification steps play crucial roles in models producing correct solutions. Building upon this observation, we also investigate which reasoning behaviors distinguish correct from incorrect solutions. As illustrated in Figure 3(d), correct solutions exhibit increased planning behaviors, potentially explaining why exploration behaviors diminish: well-structured planning facilitates clearly defined solution paths, reducing the need for exploratory detours. Additionally, correct solutions engage in verification behaviors more frequently, which slightly reduces the need for extensive analysis, as models are more confident in their reasoning. However, stronger models rely less on explicit verification overall due to robust upfront analysis and planning, which internalize many checks and reduce the need for post-hoc verification.
Based on these insights, an important direction for future research is to optimize how models allocate reasoning effort across different cognitive behaviors. Specifically, guiding models to focus on structured planning and detailed analysis, while limiting unnecessary exploration, can greatly improve solution accuracy and efficiency. Additionally, developing mechanisms that help models internalize verification during planning and analysis phases could reduce their dependence on explicit, post-hoc verification steps. Finally, creating automated techniques to dynamically adjust reasoning budgets and behaviors according to problem-specific factors like difficulty and algorithms may further boost the effectiveness of reasoning models in solving complex competitive programming tasks.
@@ 5.3 Error Patterns in Model-Generated Code Submissions
As models exhibit stronger reasoning capabilities, their solutions show substantial reductions in failure types of time limit, memory limit, and compilation errors. However, runtime errors, although somewhat reduced, do not experience as pronounced a decline, highlighting persistent challenges in edge-case handling and execution robustness. We hypothesize that one possible reason top-performing models still exhibit relatively high runtime error rates could be their tendency to pursue more aggressive and optimized coding patterns, such as employing custom data structures, in-place transformations, and pointer arithmetic. These advanced techniques, while algorithmically sound, might inherently increase the potential for execution faults [^6] For instance, a simple algorithm like summing elements of an array becomes significantly more complex when highly optimized for memory access patterns using techniques such as loop unrolling and pragma directives in C++ (e.g., #pragma omp simd, #pragma unroll)., especially in edge scenarios. Interestingly, GPT-OSS-20B displays compilation error rates comparable to weaker, non-reasoning-intensive models. We attribute this unexpected result to its cautious approach: the model often declines to generate solutions when it anticipates insufficient reasoning time, thereby triggering compilation-related failures. These findings highlight a limitation in the reinforcement learning approaches employed by current models (deepseek-r1; yang2025qwen3technicalreport), which predominantly use solution correctness as the sole reward, neglecting efficiency and memory management. Future training techniques could incorporate fine-grained reward signals targeting these attributes, enabling models to optimize not only for correctness but also for reliable and efficient code execution.
5.4 Data Contamination Study
5.4.1 No evidence of temporal performance degradation.
In Figure A2 and Appendix E.1, we evaluate contamination risk by examining whether model performance correlates with problem release dates, particularly around each model’s training cutoff. Our analysis shows no meaningful relationship between publication time and performance: models do not perform better on older tasks, nor do they exhibit any noticeable drop in accuracy on problems released after their knowledge cutoff. This lack of temporal correlation suggests that broad data leakage is unlikely and provides evidence that our benchmark is not driven by memorization effects tied to publication dates.
5.4.2 Familiarity with task statements and official solutions
Following MLE-Bench (chan2024mle-bench) on detecting data contamination, we investigate whether the models’ familiarity with task statements and official solutions affects their performance. Intuitively, if a model is more familiar with a task, it tends to perform better. Specifically, we define a model’s familiarity with a document as the mean probability it assigns to each token, conditioned on all preceding tokens. Using GPT-OSS-120B, we compute this familiarity metric and plotted it against various performance indicators, including pass rate (Figure A6) and relative score (Figure 5).
Our analysis reveals a near-zero correlation between GPT-OSS-120B’s familiarity with either task statements or official solutions and its performance. If a model’s performance strongly correlated with higher familiarity scores, this would suggest potential memorization or overfitting of training data. However, our investigation demonstrates that familiarity with task statements does not significantly predict model performance, highlighting the minimal data contamination within our benchmark. If substantial template-based contamination exists within our benchmark, one would expect higher model familiarity with the official solutions to correlate strongly with improved performance, as the model would likely have encountered identical or similarly structured solutions during training.
Figure 5: No significant positive correlation is observed between GPT-OSS-120B’s familiarity with task statements and solutions (normalized via min-max scaling) and its performance, indicating that higher familiarity does not necessarily translate to better outcomes.
5.4.3 Low similarity between models’ solutions and official solutions
Additionally, to investigate whether models rely primarily on memorization of template solutions rather than genuine reasoning, we employ the source code plagiarism detection tool Dolos (maertens2024dolos) to measure the similarity between accepted solutions generated by GPT-5 and Grok-4-Fast-Reasoning and the official solutions. GPT-5 achieves a median similarity of 0.11, while Grok-4-Fast-Reasoning has a median similarity of 0.12. In Figure A7, the majority of these solutions exhibit similarity scores below 0.3, strongly indicating that the models do not simply reproduce memorized templates. Instead, the models appear to actively reason and generate novel solutions to solve each problem. In addition, we also find negligible correlation between template similarity and model performance, suggesting minimal template contamination (Appendix E.4 and Appendix E.7).
Based on these analyses, we believe that both direct data contamination and template-based contamination are minimized in our benchmark. Moreover, to continuously mitigate potential contamination risks, our benchmark will be regularly updated every 3-6 months with newly released official contests.
5.5 Additional Analyses
Our Appendix provides additional analyses to complement the main results. Specifically, GPT-OSS-120B consistently performs best when generating solutions in C++, reflecting the language’s inherent computational efficiency advantage (Appendix E.3). Further, our detailed examination of algorithmic tags indicates that increasing reasoning-token budgets yields greater improvements on complex tasks compared to parameter scaling or training strategy variations (Appendix E.5). Finally, we find minimal variance in model performance across different prompt variants, underscoring the model’s robustness to instruction wording (Appendix E.6).
@@ 6 Conclusion
In this work, we propose LiveOIBench, a comprehensive competitive coding benchmark featuring expert-curated OI-style tasks with detailed subtask rubrics, direct comparisons to human contestant performance, continuous updates with new Olympiad tasks to prevent contamination, and an offline evaluation system ensuring accessible and reproducible assessments. We extensively evaluate $34$ models including both proprietary and open-weight models. Our results highlight that proprietary models, particularly GPT-5, achieve impressive results but fall short of top human contestants, who typically place above the 90th percentile. Among open-weight models, GPT-OSS, Seed-OSS, and Qwen-3-32B demonstrate significant progress, with GPT-OSS-120B notably narrowing the performance gap to proprietary alternatives.
Further analyses reveal that current models particularly struggle with advanced algorithmic tasks, such as dynamic programming. Additionally, our reasoning trace analysis indicates that robust model performance relies on strategically allocating exploratory and analytical reasoning behaviors. Lastly, we find stronger models reduce common failures yet persistently face runtime errors due to optimized coding techniques, suggesting refined training for efficiency and memory management. Moving forward, we envision leveraging this benchmark to further investigate inference-time scaling strategies and training methods, particularly for challenging reasoning tasks. By offering a rigorous, reproducible, and continuously updated evaluation benchmark, LiveOIBench aims to drive significant advancements in the reasoning and coding capabilities of LLMs.
@@ A.1 Task Collection
Contest Information Extraction: We developed a dedicated web crawler for each competition to extract task information directly from its official website. This includes task statements, test cases, reference and unofficial solutions, code attachments, time and memory limits, and detailed subtask specifications. We also parsed the contestant results pages and reformatted them into standardized CSV files. To do this, we copied the raw webpage content into Gemini‑2.5‑Pro and prompted it to generate CSVs with normalized headers. Each file captures contestant names, countries, total and per‑task scores, and awarded medals. After manual verification against the official data, we integrated the processed results into our contestant database. After integrating contestant results into our database, we determined medal thresholds as follows. For general contests, the Gold, Silver, and Bronze thresholds are defined by the lowest total scores among participants who received each respective medal. In contrast, for the USACO Bronze, Silver, and Gold contests, thresholds correspond to the minimum scores required to advance to the next competition level. In the USACO Platinum contest, medal assignments are based on the following score thresholds: a Gold medal is awarded to scores at or above the lowest top-ranked human score; scores above 350 receive a Silver medal; scores between 250 and 350 receive a Bronze medal; and scores of 250 or below receive no medal.
Markdown Conversion Quality Check: We sampled 40 tasks (around 10% of task statements) maintaining the same distribution across competition types as the full dataset. Our analysis revealed 4 tasks (10%) with conversion errors, primarily in nested table parsing within example inputs/outputs. While such formatting inconsistencies might impact human understanding, they have minimal effect on LLM comprehension as the core question semantic remains intact. GPT-5 fully solved three of the four tasks and passed over 70% of test cases on the remaining one, demonstrating that models are less sensitive to these formatting issues. An additional 14 tasks (35%) exhibited minor formatting issues, including header-level inconsistencies (e.g., "Input" formatted as # Input while "Output" is formatted as ### Output despite being at the same hierarchical level), incorrect superscript rendering, and uppercase/lowercase mismatches. These are inconsequential formatting artifacts that do not alter the semantic meaning of the problems. Our conversion pipeline uses Marker[1] for PDF-to-markdown conversion followed by Gemini-2.0-Flash for automated verification and correction. After manually reviewing this 40-task sample, we confirmed that the parsed task statements are indeed of high quality, with only minor formatting errors that do not impact model performance. Confident in the robustness of these results, we proceeded with batch processing the remaining tasks. We believe these slight formatting inconsistencies are inconsequential and do not compromise the validity of our evaluation.
@@ A.4 Codeforces Ratings Collection
Model Ratings: To benchmark model performance on our dataset, we calculate a corresponding Codeforces rating for each model on every contest. For each task, we present the full problem statement to the models and prompted them to generate code solutions. Using the provided subtask and test-case data, we compute the total score of each model’s solution. Once total scores were obtained, we derive Codeforces ratings for the models using the CodeElo formula (quan2025codeelobenchmarkingcompetitionlevelcode) given below, where $m$ is the expected rank of a contestant (or model) with rating $r$, compared to $n$ contestants with known Codeforces ratings $r_{(i)}$.
@@
Codeforces Profile Quality Check:
To verify the accuracy of matching contestants to their Codeforces profiles, we performed a detailed manual error analysis. We randomly sampled five contest results, selecting ten contestants from each contest. Using the same search query (contestant name, country, and "Codeforces"), we located each contestant’s Codeforces profile and verified the match by comparing the name, country, username, and yearly rating history against our stored records. Our analysis confirmed that all 50 contestant profiles matched perfectly, demonstrating 100% accuracy and confirming the high precision of our profile-matching pipeline. Nevertheless, we recognize that some contestants may lack Codeforces profiles or use alternative names on their profiles, potentially causing missed matches. Given the high precision of our pipeline, the individualized nature of ELO rating computations, and our exclusion of contests containing fewer than 15 validated contestant profiles, we anticipate that such omissions will minimally affect the overall accuracy of the model’s rating estimations.
Competitions Total Tasks Total Contests Avg. Subtasks Test Cases/Task Token Count Difficulty
@@
Table A4: Statistics of different competitions. USACO doesn’t provide subtask information.
@@
Contest Contestants Median Rating
@@ Appendix B Model Information
  • Proprietary LLMs: This category includes high-performing proprietary models such as Gemini-2.5 (comanici2025gemini), GPT-o3-Mini-High (openai2025o34), and GPT-4.1 (openai2025gpt41).
@@
  • Open-weight Thinking LLMs: These are openly available models that are empowered with inherent thinking or reasoning capabilities. This group includes Qwen3 (yang2025qwen3technicalreport) and DeepSeek-R1 (deepseek-r1), as well as those distilled from DeepSeek-R1.
@@
  • Open-weight Non-Thinking LLMs: This category consists of openly available models that are not equipped with intrinsic thinking mechanisms. This includes DeepSeek Coder-V2 (DBLP:journals/corr/abs-2406-11931), DeepSeek-V3 (DBLP:journals/corr/abs-2412-19437), Qwen2.5 (DBLP:journals/corr/abs-2412-15115), Qwen2.5-Coder (DBLP:journals/corr/abs-2409-12186), Qwen3 (DBLP:journals/corr/abs-2505-09388), Mistral (DBLP:journals/corr/abs-2310-06825) and Llama-3 (DBLP:journals/corr/abs-2407-21783).
@@
  • Refer to Table A7 and Table A8 for more details. We use the default parameter setting if it’s not listed.
  • We use the prompt at Table B for all models.
@@
Table A7: Model list of Non-Thinking LLMs with model providers and decoding settings
Non-Thinking LLMs Model Provider Temperature Top-p
GPT-4.1 (openai2025gpt41) OpenAI 1.0 1.0
Qwen2.5-72B-Instruct (DBLP:journals/corr/abs-2412-15115) Alibaba 0.7 0.8
Qwen2.5-Coder-32B-Instruct (DBLP:journals/corr/abs-2409-12186) Alibaba 0.7 0.8
Qwen2.5-Coder-14B-Instruct (DBLP:journals/corr/abs-2409-12186) Alibaba 0.7 0.8
Qwen2.5-Coder-7B-Instruct (DBLP:journals/corr/abs-2409-12186) Alibaba 0.7 0.8
Mistral-Large-Instruct-2411 (DBLP:journals/corr/abs-2310-06825) Mistral 1.0 1.0
Mistral-Small-3.1-24B-2503 (DBLP:journals/corr/abs-2310-06825) Mistral 1.0 1.0
LLaMa-4-Scout (MetaAI2025LLaMA4Multimodal) Meta 0.6 0.9
LLaMa-3.3-70B-Instruct (DBLP:journals/corr/abs-2407-21783) Meta 0.6 0.9
LLaMa-3.1-8B-Instruct (DBLP:journals/corr/abs-2407-21783) Meta 0.6 0.9
DeepSeek-V3 (DBLP:journals/corr/abs-2412-19437) DeepSeek 1.0 1.0
DeepSeek-Coder-V2-Lite-Instruct (DBLP:journals/corr/abs-2406-11931) DeepSeek 0.3 0.95
Codestral-22B-v0.1 (mistral2024codestral) Mistral 0.0 1.0
@@
Table A8: Model list with categories, including model names, organizations, reasoning budget, and decoding settings.
@@
Thinking LLMs Model Provider Reasoning Budget (Max Tokens) Temperature Top-p
GPT-5 (openai2025gpt5) OpenAI Medium (100K) - -
GPT-O3-Mini-High (openai2025o34) OpenAI High (100K) - -
GPT-OSS-120B-High (openai2025gptoss120bgptoss20bmodel) OpenAI High (128K) 1.0 1.0
GPT-OSS-20B-High (openai2025gptoss120bgptoss20bmodel) OpenAI High (128K) 1.0 1.0
GPT-OSS-120B (openai2025gptoss120bgptoss20bmodel) OpenAI Medium (128K) 1.0 1.0
GPT-OSS-20B (openai2025gptoss120bgptoss20bmodel) OpenAI Medium (128K) 1.0 1.0
Grok-4-Fast-Reasoning (xai2025grok4fast) xAI 100K 1.0 1.0
Claude-Sonnet-4.5 (anthropic2025claudesonnet45) Anthropic 120K 1.0 1.0
SEED-OSS (seed2025seedoss36binstruct) ByteDance Unlimited (128K) 1.1 0.95
Qwen3-32B (DBLP:journals/corr/abs-2505-09388) Alibaba 38K 0.6 0.95
Qwen3-14B (DBLP:journals/corr/abs-2505-09388) Alibaba 38K 0.6 0.95
QwQ-32B (qwen2025qwq32b) Alibaba 32K 0.6 0.95
Qwen3-30B (DBLP:journals/corr/abs-2505-09388) Alibaba 38K 0.6 0.95
Qwen3-8B (DBLP:journals/corr/abs-2505-09388) Alibaba 38K 0.6 0.95
Qwen3-4B (DBLP:journals/corr/abs-2505-09388) Alibaba 38K 0.6 0.95
Gemini-2.5-Pro-exp-03-25 (comanici2025gemini) Google 64K 1.0 0.95
Gemini-2.5-Flash-preview-04-17 (comanici2025gemini) Google 64K 1.0 0.95
DeepSeek-R1-01-28 (deepseek-r1) DeepSeek 32K 0.6 0.95
DeepSeek-R1-Distill-Llama-70B (deepseek-r1) DeepSeek 32K 0.6 0.95
DeepSeek-R1-Distill-Qwen-32B (deepseek-r1) DeepSeek 32K 0.6 0.95
DeepSeek-R1-Distill-Qwen-14B (deepseek-r1) DeepSeek 32K 0.6 0.95
DeepSeek-R1-Distill-Llama-8B (deepseek-r1) DeepSeek 32K 0.6 0.95
Prompt Setup Below are the instructions we use to evaluate the models, and we also include this prompt template in our appendix.
Given a competition problem below, write a solution in C++ that solves all the subtasks. Make sure to wrap your code in ’“‘<task>.cpp’ and ’“‘’ Markdown delimiters.
[BEGIN PROBLEM] <task statement> [END PROBLEM]
Time limit: <time> seconds Memory limit: <Memory> MB
Generate a solution in C++ that solves the task. Make sure to wrap your code in ’“‘<task>.cpp’ and ’“‘’ Markdown delimiters.
Prompt configuration used in all experiments.